Sol Summary – November 4th


Summary Title: The Greenhab Comes Alive
Author’s name: Guy Murphy
Mission Status: Active.

Sol Activity Summary: A long morning EVA was undertaken to collect samples as part of the micro-meteorite samples. The Greenhab was partially planted up yesterday and routine watering commenced today. 3 cooked meals were prepared today, including a special dinner to mark Melbourne Cup. Information was collected on radio inventories and the condition of several space suits at the request of Mission support.

Look Ahead Plan: We are planning the remaining week tightly to complete our planned sample collection and allocate time to complete end-of mission tasks.

Anomalies in work: None

Weather: Another clear sunny day on Mars.

Crew Physical Status: Crew continue to be in excellent health.

EVA: See EVA report.

Reports to be filed: Sol Summary, EVA Report, Journalists Report, Science Report, Operations Reports, Photos.

Support Requested: Special delivery of supplies were greatly appreciated, especially the milk. There are 3 bags of rubbish to be collected from the back air lock (2 burnable, 1 non-burnable).

Operations Report – November 4th

Crew 214 Operations Report 4-11-2019
SOL: 8
Name of person filing report: Andrew Wheeler
Non-nominal systems: Opportunity, Suit 10
Notes on non-nominal systems: Opportunity brakes do not work, Suit 10 is not functioning

Generator: run
Hours run: 13h 15m
From what time last night: 1845
To what time this morning: 0800
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night): 94%

Diesel Reading – 60%
Station Propane Reading – 82%
Ethanol Free Gasoline: N/A

Water (loft tank): 32 gallons
Water Meter: 1455633 units
Water (static tank): 246 gallons
Static to Loft Pump used – no
Water in Green Hab: 225 gallons
Water in Science Dome: 0 gallons
Toilet tank emptied: yes

Deimos rover used: not at hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –

Spirit rover used: yes
Hours: 107.4
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 52%
Currently charging: yes

Opportunity rover used: no
Hours: 65.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Curiosity rover used: no
Hours: 113.4
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 100%
Currently charging: yes

Notes on rovers: no

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used: 0 Gals
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director

HabCar used and why, where? N/A

CrewCar used and why, where? N/A

General notes and comments: N/A

Summary of internet: Functional
Summary of suits and radios: 5 Radios and 3 earpieces that need servicing have been identified and set aside, the battery for suit 10 is fully charged and it is Suit 10 that is not functional, further information on Suit 5’s helmet has been forwarded to Capcom.
Summary of Hab operations: All systems nominal
Summary of GreenHab operations: Atmospheric equalization to the green hab was carried out during the day, plants were watered.
Summary of ScienceDome operations: All systems nominal. Sample processing following EVA was carried out.
Summary of RAM operations: Not used
Summary of any observatory issues: N/A
Summary of health and safety issues: All crew healthy
Questions, concerns and requests to Mission Support: Acknowledge that the request for 6 radio batteries is being processed.

Journalist Report – November 4th

Mon 4 Nov Sol 8
by Sandy Dance

The two science experiments proceed apace. Dianne’s work happens at each meal, busy weighing and analyzing the food data. Meanwhile, this morning Andrew and I went on an EVA back to the micrometeorite grid, and collected 28 squares worth of particles. We got through more squares today because we had more time, and just went to the site and did one thing without diversions. We may also just be getting more skillful at the complex process of waving the magnet, carefully bagging the collected particles, and attaching the next bag onto the magnet. Very satisfying, but nevertheless, after our return I was quite exhausted.

Tonight the Melbourne Cup runs at 9pm, so we have been preparing special food: jambalaya, sauteed vegetables, and mashed potatoes. To be followed with freshly cooked brownies and a bottle of sparkling, not wine, but Perrier mineral water. Oooh, fancy! Whether have the stamina to stay up for the actual horse race is as yet unknown.

EVA Report – November 4th

Crew 214 EVA Report 04-11-2019
EVA # 7
Author: Andrew Wheeler
Purpose of EVA: Geological Sampling

Start time: 0945
End time: 1215

Narrative: The fifth EVA for science saw us returning to the Tank Wash location #2 at 518477E 4253767N where we continued sampling the 1m x 1m plots) for micrometeorites. The longer EVA allowed us to collect 28 samples. We have now completed 40% of the grid and anticipate 3 x normal length or 2 x longer duration EVAs will complete this sampling program.

Destination: Drove from the hab along Cow Dung Road to Tank Wash and then returned to the hab.

Coordinates: 518477E 4253767N
Participants: Andrew Wheeler, Sandy Dance
Road(s) and routes per MDRS Map: Cow Dung Road to Tank Wash
Mode of travel: Spirit Rover

Science Report – November 4th

Crew 214 Science Report 4 November 2019
SOL: 8
Crew 214 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler

Science Operations: The fifth scientific EVA for the mission saw us returning to Tank Wash #2 and collecting 28 additional samples for the micrometeorite search. This is 40% of the planned samples.

Green Hab: Following the commencement of planting within the green hab, access through the tunnel between the main hab and the green hab has been opened. The plants were watered with approximately 3.5 litres of water, with a remaining 4.5 litres remaining in the watering can. The plant soil appeared fairly moist, so it is suggested crew refrain from over watering the new plants.

Science Dome: Samples returned from the morning’s EVA have now been processed within the science dome and limited observations of material recovered through the binocular microscope have been carried out.

Closed Loop Food Waste Study: The day’s waste for Sol 8 is: 51 grams; consisting of 159 calories, 3g fat, >1g saturated fat, 199mg sodium, 42g carbohydrates, 3g fibre, 13g sugar, 6g protein. Extra food supplies were received and added to the baseline stock database.

EVA: Preparations for an EVA to White Moon at the junction between Brahe Highway and Cow Dung Road have been carried out.

EVA Suit Maintenance: Images of helmet 5 have been forwarded to CapCom detailing the partial failure, the battery to Suit 10 has been examined and it is the suit itself that is faulty.

Additional Activities: Nil.

Supplement Operations Report – November 4th

MDRS Supplemental Operations Report 4-NOV-2019

Name of person filing report: Shannon Rupert

Reason for Report: Briefing on Work Party held at MDRS from Oct 22-25, 2019

Non-nominal systems: n/a

Action taken for non-nominal systems: n/a

Generator check, note if oil and coolant added: n/a

ScienceDome AC unit, on and operational: Needs refrigerant, but otherwise is ready to go. Will contact provider.

Solar— Ran nominally and without generator backup throughout the work party

Diesel Reading – % not checked

Propane Reading – 40 % (since then it has been refilled)

Ethanol Free Gasoline – n/a gallons.

Water (director’s tank) – n/a gallons

Water (static tank) – n/a gallons. Static tank was cleaned during
the work party.

Water (GreenHab) – n/a gallons

Water (loft) – This tank still needs to be cleaned and will be cleaned during the crew changeover from Crew 214 to 215. We also need to replace the current hose with a 4-6 foot RV hose (to be ordered)

Static to Loft Pump used – yes

Water Meter: n/a

Toilet tank emptied: yes. We also filled and emptied the tank three times to flush it after the summer break.

Deimos rover used: Still not operational. We are in the process of getting a trailer to take it for repairs.


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: Yes, brought back to MDRS from town. Operation nominal.


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Yes, brought back to MDRS from town. Operation nominal


Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: Yes, brought back to MDRS from town.
Operation nominal, however brakes are weak to not working. We will change the brake fluid and bleed the brakes again (3rd time), if this does not work we will need to take it for repairs.


Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: Yes, brought back to MDRS from town. Operation nominal.


Beginning charge:

Ending charge:

Currently charging:

Notes and action taken on rovers: Rovers need their names redone. I’m thinking about painting them instead of using stickers again.

Monthly check and fill of rover batteries (first workday of the month)
Batteries were filled with 20 gallons of distilled water this summer.
They will need to be rechecked soon.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3). Yes, moved to their place by the Outpost for the season. The 300 was in town, and still is, as it will not start. We will need to go to town and tow it back here. One of the 350’s would not start so we need to troubleshoot

Reason for use: see above

Oil Added? No, all were fine.

ATV Fuel Used: 5 Gals

# Hours the ATVs were Used: not recorded

Notes and action taken on ATVs: see above

HabCar running? If used, why, where? Yes, but needed to be jumped when we first got here. Used to haul water.

CrewCar running, If used why, where? Yes, used to transport crew and supplies back and forth to town and Grand Junction. A sensor went out, and one of the volunteers was able to replace it. The check engine light is still on though, as they could not get it to turn off by disconnecting the battery.

Campus wide inspection, if action taken, what and why? Vegetation was removed from around the Hab.

Evidence of rodents, where and action taken: Hab and both trailers showed signs of rodent (packrat) occupation. Two packrats were relocated to Cowboy Corner, one from the intern trailer, the other from the Hab.

Summary of internet: Hab and Astronomy accounts nominal but business account isn’t working. Suspect router damage. There is supposed to be a technician here today (although they were scheduled for last week, so I’m not going to hold my breath).

Summary of suits and radios: Suits were delivered by NorCal on October 25 and they are in beautiful shape. The prototype suit has been fitted with a video camera and will be used for Mars VR work. This gives us 6 two piece suits, which are located on new cabinets in
the EVA room and 4 one piece suits, which are located on the lower deck in the main room. One of the two piece suits has been fitted with a 3D printed neck ring, which is about 1/3 the weight of the traditional neck rings. NorCal is interested in seeing how it holds
up. The two piece suits have also been fitted with new fans, with adjustable air flow. The fans are new in the one piece suits as well, but not adjustable.

Radios were not checked during the work party, but Crew 214 has found them to be in poor shape. Once internet is restored we will get an accounting of them and order what we need.

Summary of Hab operations: The front porch was repaired. The static water tank was cleaned. The wall heater downstairs was removed. The lower deck was refitted with IKEA cabinetry. While not the cubbies for the suits we were hoping for, it fills the lower deck and is
functional. They also gave us a new bench for the bathroom. The painting did not get done before the installation so that will happen during the May work party. The upper deck received a completely new kitchen. The stove was not installed as it was electric. We will go
pick up a new gas stove from IKEA when they get it. The countertop, which is quartz, was measured and ordered and should arrive soon. A new two basin sink was installed on the temporary counter and new plumbing was installed, so both cold and hot water are running and there is no more leak in the drain. The kitchen also has a new microwave and refrigerator and it several feet longer than the old one. In addition, the upper deck has new chairs for the table and two new chairs that replaced the loungers. Unfortunately the two leather
chairs are not functional as loungers so we will be replacing them with some kind of lounger that fits in the new space. We will try to use the old loveseat pieces but the parts for those were disposed of during the refit, so we may not be able to do so. Lights, chargers and
additional mattresses were put into the staterooms. The old blue mattresses were removed. Water heater still needs to be removed and replaced with a new unit. Some new kitchen ware was ordered as some was deemed no longer in good enough condition for use.

Summary of director’s house operations: Nothing to report.

Summary of GreenHab operations: IKEA refitted the GreenHab with some cabinets and a couple of new benches and gave us some planters. The benches need to be coated with a water seal and the cabinets are metal, so care must be taken not to corrode them. I will make up a diagram of the new space soon.

Summary of ScienceDome operations: Nothing was done in the ScienceDome. We attempted to put in the center island, but parts were missing (it was nice to know that even IKEA can have missing parts!) so this refit is on hold until the May work party. A new heating/ac
units was mounted on the power wall over the summer, but is still not fully operational. It was installed to cool the batteries in summer, but can also be used as a heater if needed.

Summary of RAM operations: Nothing to report. David is working on a plan to fix the low amp breakers than trip with the rovers.

Summary of any observatory issues: Nothing to report. Peter needed David and Atila’s help to calibrate the focus on the robotic scope but all is well now.

Summary of health and safety issues: Fire extinguishers need their annual inspection

List and summary of projects: see above. Also, the new intern trailer was moved into a space next to mine in August. We are working to get all systems functional on the trailer before the bad weather hits. Water system has required small repairs and the propane will be
installed this Thursday. The best thing about this trailer (other than it is a four season trailer) is that its solar system also has an inverter so it is currently operating under it’s own generated power. This was a surprise for me as my solar only operates the 12V system
and not the 110 power. I may have to move! Just kidding.

Questions, concerns, supplies needed and requests: We need some general replacement supplies and I will be ordering them soon. We currently have the flatbed trailer loaded with materials needing to go to the dump. I will need to find a way to get it there.

Copyright © The Mars Society. All rights reserved. | Main Site